Maria can shovel snow from her driveway in 60 minutes. Jim can do the same job in 55 minutes. How long would it take maria and jim to shovel the driveway if they worked together?
Thank you

Answer by Alan3354(69443)

M does 1/60 of the job per minute.
J does 1/55 per minute.
Together, they do 1/60 + 1/55 per minute
= 60/3300 + 55/3300
= 115/3300 per minute
--> 3300/115 minutes per job
= 660/23 minutes
=~ 28.7 minutes
-----------------
Use the shortcut, product/sum
60*55/(60 + 55)
= same answer